Solve the system of equations and give value for x:
-5x = -10 - 5y
3x + 2y =16
Please help with this. Thanks.

Answer by stanbon(75887)

-5x = -10 - 5y
3x + 2y =16
----
Rearrange:
5x - 5y = 10
3x + 2y = 16
----
Modify for elimination:
15x - 15y = 30
15x + 10y = 80
---------------------
Subtract and solve for "y":
25y = 50
y = 2
-----
Solve for "x":
x - y = 2
x - 2 = 2
x = 4
-----------------
Cheers,
Stan H.
=================